OpenStudy (anonymous):

how do you put x - y = -1 in slope intercept form

OpenStudy (austinl):

Slope intercept form is \(\color{blue}{y=mx+b}\) How would you get it to look like that?

OpenStudy (anonymous):

idk

OpenStudy (austinl):

Well we need to get y alone on the left side of the equation, what would you do to achieve that?

OpenStudy (anonymous):

subtract y from both sides???

OpenStudy (austinl):

I would subtract x from both sides first. Try that and tell me what you get :)

OpenStudy (anonymous):

-x-y=1??

OpenStudy (austinl):

We have, \(x-y=-1\) Right? Subtract x from both sides. \((\cancel{x}-y)-\cancel{x}=-1-x\) \(-y=-x-1\)

OpenStudy (austinl):

Now we need y not -y, so what would you do to both sides to do that? (Hint: this is easy)

OpenStudy (anonymous):

divide by -y ?

OpenStudy (austinl):

Nope, negate both sides. Divide each side by negative 1.

OpenStudy (anonymous):

how?

OpenStudy (austinl):

\(\dfrac{-y}{-1}=\dfrac{-1-x}{-1}\)

OpenStudy (anonymous):

what should i get?

OpenStudy (austinl):

what is a number divided by one?

OpenStudy (anonymous):

1

OpenStudy (austinl):

No, any number divided by one? What is it.

OpenStudy (anonymous):

that number

OpenStudy (austinl):

Exactly, and what if that one is negative?

OpenStudy (anonymous):

ooo i got it is the answer y= 1-x

OpenStudy (anonymous):

????

OpenStudy (austinl):

\(\dfrac{-x}{-1}=~?\)

OpenStudy (anonymous):

yes

OpenStudy (austinl):

No, what does it equal?